SMALL SULFUR CLUSTERS IN THE GAS PHASE: PURE ROTATIONAL SPECTROSCOPY OF S_{3} (THIO-ZONE) AND S_{4}

\underline{S. THORWIRTH}, M. C. McCARTHY, C. A. GOTTLIEB, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A.; H. GUPTA, J. F. STANTON, University of Texas at Austin, Department of Chemistry and Biochemistry, Institute for Theoretical Chemistry, Austin, TX 78712, U.S.A..

The C_{2v} forms of thiozone, S_{3}, and tetrasulfur, S_{4}, have been investigated by Fourier transform microwave spectroscopy of a supersonic molecular beam. Both molecules represent the global minima on their respective potential energy surfaces. Besides the parent molecules the mono-^{34}S isotopomers have been investigated in natural abundance. For thiozone, a substantial number of vibrational satellites has been detected in addition. The analysis of the spectroscopic data will be discussed and a comparison to recent high-level quantum chemical calculations at the CCSD(T) level of theory will be given. From a calculation of the cubic force field the rotation-vibration interaction constants have been calculated and the mixed experimental-theoretical r_{e} structures have been obtained.